A coffee merchant combines coffee costing $5.00 per pound with coffee costing $3.50 per pound. How many pounds of each should be used to make 30 lb of a blend costing $4.20 per pound?
Answer by Alan3354(69443)

Here's a similar problem:
-------------
how many liters of a 70% acid solution must be mixed with a 15% acid solution to get 385L of a 60% acid solution?
-------------------
s = amount of 70%
f = amount of 15%
-------
s + f = 385 (total solution)
0.7s + 0.15f = 0.6*385 (total acid)
-------
f = 385 - s
Sub for f in 2nd equation
0.7s + 0.15(385-s) = 231
70s + 5775 - 15s = 23100
55s = 17325
s = 315 liters of 70%
------------
f = 70 liters of 15%